Simplify
-/-3 +4/ + 5 - 7 =

Answer by ttlpkg32(2)

-|-3+4|+5-7=
- 3+4+5-7=
-7+5-7=
-2-7=-9
Absolute values are always positive. But if you have a negative sign before that absolute value sign, that changes the number number to a negative number. Good Luck. Bill